sorry i wrote the last problem wrong it should of read
Find probability
suppose one card is selected at random from an ordianry deck of 52 playing cards
let
a = event a queen is selected
b= event a diamond is selected
determine P(B/A)

Answer by Edwin McCravy(6939)


sorry i wrote the last problem wrong it should of read 
Find probability 
suppose one card is selected at random from an ordianry 
deck of 52 playing cards
let
A = event a queen is selected
B = event a diamond is selected 
determine P(B|A)

There awo two ways to find conditional probabilities

Method 1 (reduce the sample space)

You are given that you have selected a queen.
So you are given that you drew one of the four Queens.
So exactly one of those four queens you could have selected
is a diamond.

So P(B|A) = 1/4.

Use the formula:

          P(X and Y)  
P(X|Y) = ------------- 
             P(Y)

          P(B and A)  
P(B|A) = ------------- 
             P(A)

                    P(diamond and Queen)  
P(diamond|Queen) = ---------------------- 
                         P(Queen)

P(diamond and queen) = P(queen of diamonds) = 1/52

P(Queen) = 4/52 

                    1/52     1
P(diamond|Queen) = ------ = --- = 1/4 
                    4/52     4

Either way it comes out 1/4.

Edwin